OpenStudy (anonymous):

log3(3 square root 3)= ?

OpenStudy (anonymous):

\[\log_3(3\sqrt{3}) = \log_3(3 \cdot 3^{1/2}) = \log_3(3^{3/2}) = 3/2\]